# Hyper-Compression: Model Compression via Hyperfunction

### v3

@@ 5 Experiment and Analysis
1 225 0.02 99.81 7.09$\times$
2 225 0.02 97.07 7.08$\times$
3 225 0.02 90.11 7.08$\times$
1 225 0.1 99.21 7.92$\times$
2 225 0.1 96.46 7.88$\times$
3 225 0.1 99.71 7.88$\times$
1 361 0.02 99.87 6.38$\times$
2 361 0.02 97.90 6.38$\times$
3 361 0.02 99.39 6.38$\times$
1 361 0.1 99.28 7.73$\times$
2 361 0.1 99.10 7.71$\times$
3 361 0.1 99.65 7.68$\times$
1 225 / 361 0.02 99.89 7.08$\times$
2 225 / 361 0.02 98.12 7.08$\times$
3 225 / 361 0.02 99.58 7.08$\times$
1 225 / 361 0.1 99.08 7.73$\times$
2 225 / 361 0.1 98.85 7.71$\times$
3 225 / 361 0.1 99.51 7.68$\times$
Figure 9: Parameter sensitivity results for different hyper-parameters on UNet.
As for UNet and MobileNetV3, our method can compress UNet and Pruned-UNet by 7.87×7.87$\times$7.87 × and 7.05×7.05$\times$7.05 × with the performance loss contained in 1%, as shown in Table 5. Particularly, our method succeeds in combination with other model compression methods such as pruning to achieve an even higher compression ratio. In UNet, the total compression ratio is 17.74×17.74$\times$17.74 × with the performance loss $3.41\%$.
No post-hoc retraining. Table 7 compares the number of tokens different compression methods use to retrain the compressed models, which highlights that our hyper-compression is one-shot. This is because discrepancies between the original and decoded parameters in the hyper-compression are minor, at magnitudes ranging from $10^{-4}$ to $10^{-3}$. Then, the impact on error accumulation through layer-by-layer propagation is acceptable. This advantage is particularly useful in industry, where often 1) training data are inaccessible, and///or curating data is costly; 2) no computing resources are supplied for retraining.
Table 7: The data tokens used in pre-training process compared with other models. Our model compression method don’t need any pretraining or fine-tuning data.
@@
LLaMA2-7B et al. (2023) 2T Unknown
Sheared-LLaMA-1.3B Xia et al. (2024) 50B RedPajama
TinyLLaMA Zhang et al. (2024) 3T Slimpajama +StarCoder
@@
OPT Zhang et al. (2022b) 300B RoBERTa + The Pile + PushShift.io Reddit
Pythia Biderman et al. (2023) 300B The Pile
@@
OpenLLaMA v1 Geng and Liu (2023) 1T RedPaJama
OpenLLaMA v2 Geng and Liu (2023) 1T Falcon refined-web + StarCoder + Parts of RedPaJama
@@
Affordable inference time. A network is a hierarchical structure. To expedite the inference time, our approach is to leverage this hierarchy by parallelizing the decoding and inference processes. The key is to complete the decoding of a layer before using this layer to infer. As shown in Table 8, “Original” refers to the time required for a single inference of the original UNet, and “Ours” refers to the time required for one inference by employing parallel decoding and inference processes. It can be seen that the hyper-compression only increases the total inference time moderately.
Table 8: Timing results for different batch sizes on two devices using two methods with a UNet network.
@@
Short compression time. As Table 9 shows, our method can compress models very fast. This efficiency primarily stems from using matrix operations for most compression processes and the K-D tree to expedite the search. Additionally, by treating the compression tasks between layers as independent operations, we implement a parallel strategy to further decrease the compression time.
Table 9: The compression time of our method on UNet, MobileNetV3, and LLaMA2-7B.
Model Time (s)
UNet + HF 30.00
MobileNetV3 + HF 11.63
Model Time (min)
LlaMA2-7B + HF 53
Sheared-LlaMA-1.3B + HF 7
6 Ablation Study and Parameter Sensitivity
As shown in Table 10, we conduct ablation experiments based on UNet and MobileNetV3 on the aforementioned two acceleration techniques to evaluate their independent contributions in computational speed. It is seen that both k-d tree and matrix operations leads to a substantial enhancement in computational efficiency. Notably, the k-d tree technique becomes more pronounced when the matrix operations are applied; and vice versa.
Table 10: The results of ablation study on acceleration techniques.
Model K-D Matrix Compression
Tree Operations Time (min)
UNet Ronneberger et al. (2015) 231.6
176.9
54.8
0.5
MobileNetV3 Howard et al. (2019) 52.29
3.13
45.50
0.19
We use UNet as an example to conduct a parameter sensitivity test on three aforementioned hyperparameters: $M$, $U$, and $l$. $M$ means the maximum value of categories that is set for all layers of the model, $U$ means the list of the number of sample nodes in the center square, and $l$ means the length of the side of the center square. As shown in Table 9, different values of $M$, $U$, and $l$ only moderately impact the model’s performance and compression ratio, which means our algorithm is robust to hyperparameters.
@@ 7 Conclusion and Future Work
In this study, we have proposed hyper-compression, a novel and general-purpose methodology for model compression that leverages the trajectory density of some dynamic system to encode the parameters of the target network. We have conducted comprehensive and systematic experiments on various architectures, including LlaMA2-7B, UNet, and MobileNetV3, demonstrating that our method is both user-friendly and scalable. Nevertheless, it represents merely an initial foray into the whole landscape. Future directions include 1) beyond classical theory, exploring other possibilities such as implicit neural representation Chitturi et al. (2023) to strike a better balance between compression time, inference time, and compression ratio; 2) considering distributions of weights of the target network to secure a higher compression rate. We believe that hyper-compression can contribute to Moore’s law of model compression in the near future, i.e., the compression efficiency can be doubled annually, as a solution for the stagnation of hardware Moore’s law.

### v4

@@ 6 Ablation Study and Parameter Sensitivity
[…]
TABLE XIV: The data tokens used in pre-training process compared with other models. Our model compression method don’t need any pretraining or fine-tuning data.
@@
LLaMA2-7B [17] 2T Unknown
Sheared-LLaMA-1.3B [51] 50B RedPajama
TinyLLaMA [56] 3T Slimpajama
StarCoder
@@
OPT [59] 300B RoBERTa
The Pile
PushShift.io Reddit
Pythia [4] 300B The Pile
@@
OpenLLaMA v1 [21] 1T RedPaJama
OpenLLaMA v2 [21] 1T Falcon refined-web
StarCoder
Parts of RedPaJama
@@
TABLE XV: Timing results for different batch sizes on two devices using two methods with a UNet network.
@@
7 Discussion
The choice of hyperfunction for Hyper-Compression. We choose the irrational winding on the torus as our hyperfunction because of its high computational efficiency. The irrational winding as a candidate hyperfunction is extremely simple and efficient to compute: $w_{n}=\tau(\theta\cdot\alpha_{n})$, where $\tau(z)=z-\lfloor z\rfloor$, and $\alpha_{n}$ is irrational. This means that to recover weights of the original network, we only need a single multiplication and modulo operation per weight, which is instrumental in reducing the delay in inference. Meanwhile, finding $\theta$ does not need to deal with a partial differential equation (PDE). Thus, we can leverage matrix operations to compress and decompress entire layers in parallel, thereby allowing our method to scale to multi-billion parameter models, e.g., compressing LLaMA2-70B in 33.6 minutes (please see Table XI).
Our work establishes a relationship between the ergodicity of the dynamic system and model compression. Other famous ergodic dynamic systems should also have the potential of compression: i) Chaotic Maps [2]: While they can produce dense trajectories, they often exhibit sensitivity to initial conditions, rendering it hard to determine the initial point. Moreover, chaotic equations or formulas are serial, which need a long time to recover the weights, creating a delay in network inference. ii) Linear Congruential Generators (LCGs, [6]): They are also serial systems, and are hard to parallelize, facing the same issue. iii) Quasi-Monte Carlo Sequences [7]: systems with faster mixing rates could potentially achieve lower reconstruction errors for the same trajectory length. However, they typically require storing a base sequence or complex generation rules, making them hard to scale to large models.
We acknowledge that other dynamic systems might offer advantages in specific scenarios and are an exciting direction for future work. However, it can be seen that these systems face the challenge of scalability, which is the key matter in the era of large models. In contrast, the irrational winding strikes a good balance among strong scalability and competitive performance across diverse models.
Discussion of theory. Theorem 12 provides a probabilistic upper bound on the approximation error for a single linear layer. This bound is derived using concentration inequalities and the properties of irrational winding. We emphasize that this bound is
  • Asymptotic: It describes the dynamic behavior of the compression error as $N_{0}$ changes.
  • Worst-case: It holds for the worst-case input distribution (uniform on the sphere) and does not fully account for the structure in real data.
In practice, we observe that the actual errors are significantly lower than the theoretical bound. This is because we employ engineering techniques like scaling and KD-tree-based trajectory assignment that can minimize the reconstruction error by adapting the hyperfunction to the weight distribution, particularly dealing with outliers appropriately.
However, we think that our theoretical derivation is meaningful in terms of informing how the key factors determine the recovering error. To evaluate the role of $N_{0}$, we conduct experiments on synthetic weight tensors under controlled settings: We fix $\Theta=10^{8}$ and sample input data $\mathbf{X}\in[0,1]^{m\times N_{0}}$ from a uniform distribution. For each row of $\mathbf{X}$, we find the corresponding $\theta$. Four configurations are evaluated with $m\in\{128,256,512,1024\}$. For each $N_{0}$, we compute the mean L1 error across 100 independent trials. As demonstrated in Figure 12, the L1 error roughly exhibits an increasing trend with a growing $N_{0}$, which agrees with our theoretical prediction.
@@ 8 Conclusion and Future Work
In this study, we have proposed Hyper-Compression, a novel and general-purpose methodology for model compression that leverages the trajectory density of some dynamic system to encode the parameters of the target network. We have conducted comprehensive and systematic experiments on various architectures, including LlaMA series, Qwen series, ResNet series, UNet, and MobileNetV3, demonstrating that our method is both user-friendly and scalable. Nevertheless, it represents merely an initial foray into the whole landscape. Future directions include 1) developing a fused decode-compute method that integrates weight decoding directly into the matrix multiplication kernel for inference speedup; 2) beyond classical theory, exploring other possibilities such as implicit neural representation [10] to strike a better balance between compression time, inference time, and compression ratio; 3) considering distributions of weights of the target network to secure a higher compression rate. We believe that Hyper-Compression can contribute to Moore’s law of model compression in the near future, i.e., the compression efficiency can be doubled annually, as a solution for the stagnation of hardware Moore’s law.
Reference
  • [1] I. Ahmed, S. Islam, P. P. Datta, I. Kabir, N. U. R. Chowdhury, and A. Haque (2025) Qwen 2.5: a comprehensive review of the leading resource-efficient llm with potentioal to surpass all competitors. Authorea Preprints. Cited by: §1, §5, §5.
  • [2] K. T. Alligood, T. D. Sauer, J. A. Yorke, and D. Chillingworth (1998) Chaos: an introduction to dynamical systems. SIAM Review 40 (3), pp. 732–732. Cited by: §7.
  • [3] Y. An, X. Zhao, T. Yu, M. Tang, and J. Wang (2024) Fluctuation-based adaptive structured pruning for large language models. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38, pp. 10865–10873. Cited by: §2.1.
  • [4] S. Biderman, H. Schoelkopf, Q. G. Anthony, H. Bradley, K. O’Brien, E. Hallahan, M. A. Khan, S. Purohit, U. S. Prashanth, E. Raff, et al. (2023) Pythia: a suite for analyzing large language models across training and scaling. In International Conference on Machine Learning, pp. 2397–2430. Cited by: TABLE XIV.
  • [5] Y. Bisk, R. Zellers, J. Gao, Y. Choi, et al. (2020) Piqa: reasoning about physical commonsense in natural language. In Proceedings of the AAAI conference on artificial intelligence, Vol. 34, pp. 7432–7439. Cited by: §5.
  • [6] J. Boyar (1989) Inferring sequences produced by pseudo-random number generators. Journal of the ACM (JACM) 36 (1), pp. 129–141. Cited by: §7.
  • [7] R. E. Caflisch (1998) Monte carlo and quasi-monte carlo methods. Acta numerica 7, pp. 1–49. Cited by: §7.
  • [8] Y. Cai, Z. Yao, Z. Dong, A. Gholami, M. W. Mahoney, and K. Keutzer (2020) Zeroq: a novel zero shot quantization framework. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pp. 13169–13178. Cited by: §5.
  • [9] V. K. Chauhan, J. Zhou, P. Lu, S. Molaei, and D. A. Clifton (2024) A brief review of hypernetworks in deep learning. Artificial Intelligence Review 57 (9), pp. 250. Cited by: §1, §3.1.
  • [10] S. R. Chitturi, Z. Ji, A. N. Petsch, C. Peng, Z. Chen, R. Plumley, M. Dunne, S. Mardanya, S. Chowdhury, H. Chen, et al. (2023) Capturing dynamical correlations using implicit neural representations. Nature Communications 14 (1), pp. 5852. Cited by: §8.
  • [11] K. Cobbe, V. Kosaraju, M. Bavarian, M. Chen, H. Jun, L. Kaiser, M. Plappert, J. Tworek, J. Hilton, R. Nakano, et al. (2021) Training verifiers to solve math word problems. arXiv preprint arXiv:2110.14168. Cited by: §5.
  • [12] I. P. Cornfeld, S. V. Fomin, and Y. G. Sinai (2012) Ergodic theory. Vol. 245, Springer Science & Business Media. Cited by: §1, §3.2, Definition 1, Definition 2.
  • [13] M. De Marneffe, M. Simons, and J. Tonhauser (2019) The commitmentbank: investigating projection in naturally occurring discourse. In proceedings of Sinn und Bedeutung, Vol. 23, pp. 107–124. Cited by: §5.
  • [14] T. Dettmers, M. Lewis, Y. Belkada, and L. Zettlemoyer (2022) Gpt3. int8 (): 8-bit matrix multiplication for transformers at scale. Advances in neural information processing systems 35, pp. 30318–30332. Cited by: §2.1, §2.1.
  • [15] N. Ding, Y. Qin, G. Yang, F. Wei, Z. Yang, Y. Su, S. Hu, Y. Chen, C. Chan, W. Chen, et al. (2023) Parameter-efficient fine-tuning of large-scale pre-trained language models. Nature Machine Intelligence 5 (3), pp. 220–235. Cited by: §1.
  • [16] E. Dupont, A. Goliński, M. Alizadeh, Y. W. Teh, and A. Doucet (2021) Coin: compression with implicit neural representations. arXiv preprint arXiv:2103.03123. Cited by: §2.2.
  • [17] H. T. et al. (2023) Llama 2: open foundation and fine-tuned chat models. External Links: 2307.09288 Cited by: §1, TABLE III, TABLE V, TABLE V, §5, §5, TABLE XIV.
  • [18] E. Frantar and D. Alistarh (2022) Optimal brain compression: a framework for accurate post-training quantization and pruning. Advances in Neural Information Processing Systems 35, pp. 4475–4488. Cited by: §2.1.
  • [19] E. Frantar and D. Alistarh (2023) Sparsegpt: massive language models can be accurately pruned in one-shot. In International Conference on Machine Learning, pp. 10323–10337. Cited by: §2.1.
  • [20] E. Frantar, S. Ashkboos, T. Hoefler, and D. Alistarh (2022) Gptq: accurate post-training quantization for generative pre-trained transformers. arXiv preprint arXiv:2210.17323. Cited by: TABLE X, TABLE X, TABLE VI, TABLE IX, TABLE IX, §5.
  • [21] X. Geng and H. Liu (2023) Openllama: an open reproduction of llama. URL: https://github. com/openlm-research/open_llama. Cited by: TABLE XIV, TABLE XIV.
  • [22] A. Ghorbel, W. Hamidouche, and L. Morin (2024) NeRV++: an enhanced implicit neural video representation. In 2024 IEEE International Conference on Visual Communications and Image Processing (VCIP), pp. 1–5. Cited by: §2.2.
  • [23] C. Guo, Y. Qiu, J. Leng, X. Gao, C. Zhang, Y. Liu, F. Yang, Y. Zhu, and M. Guo (2022) Squant: on-the-fly data-free quantization via diagonal hessian approximation. arXiv preprint arXiv:2202.07471. Cited by: §5.
  • [24] D. Ha, A. M. Dai, and Q. V. Le (2017) HyperNetworks. In International Conference on Learning Representations, Cited by: §2.3.
  • [25] K. He, X. Zhang, S. Ren, and J. Sun (2016) Deep residual learning for image recognition. In Proceedings of the IEEE conference on computer vision and pattern recognition, pp. 770–778. Cited by: §1, TABLE III, TABLE III, §5, §5.
  • [26] A. Howard, M. Sandler, G. Chu, L. Chen, B. Chen, M. Tan, W. Wang, Y. Zhu, R. Pang, V. Vasudevan, Q. V. Le, and H. Adam (2019) Searching for mobilenetv3. External Links: 1905.02244 Cited by: §1, TABLE III, TABLE VIII, §5, TABLE XII.
  • [27] E. J. Hu, P. Wallis, Z. Allen-Zhu, Y. Li, S. Wang, L. Wang, W. Chen, et al. (2022) LoRA: low-rank adaptation of large language models. In International Conference on Learning Representations, Cited by: §2.1.
  • [28] W. Huang, Y. Liu, H. Qin, Y. Li, S. Zhang, X. Liu, M. Magno, and X. Qi (2024) Billm: pushing the limit of post-training quantization for llms. arXiv preprint arXiv:2402.04291. Cited by: §2.1.
  • [29] A. Katok, A. Katok, and B. Hasselblatt (1995) Introduction to the modern theory of dynamical systems. Cambridge university press. Cited by: §1, Theorem 3.
  • [30] D. Khashabi, S. Chaturvedi, M. Roth, S. Upadhyay, and D. Roth (2018) Looking beyond the surface: a challenge set for reading comprehension over multiple sentences. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), pp. 252–262. Cited by: §5.
  • [31] A. Krizhevsky, I. Sutskever, and G. E. Hinton (2012) Imagenet classification with deep convolutional neural networks. Advances in neural information processing systems 25. Cited by: §5.
  • [32] H. J. Levesque, E. Davis, and L. Morgenstern (2012) The winograd schema challenge.. KR 2012 (13th), pp. 3. Cited by: §5.
  • [33] J. Lin, J. Tang, H. Tang, S. Yang, W. Chen, W. Wang, G. Xiao, X. Dang, C. Gan, and S. Han (2024) AWQ: activation-aware weight quantization for on-device llm compression and acceleration. Proceedings of Machine Learning and Systems 6, pp. 87–100. Cited by: §2.1, TABLE X, TABLE IX, §5, §5.
  • [34] Z. Liu, B. Oguz, C. Zhao, E. Chang, P. Stock, Y. Mehdad, Y. Shi, R. Krishnamoorthi, and V. Chandra (2023) Llm-qat: data-free quantization aware training for large language models. arXiv preprint arXiv:2305.17888. Cited by: §2.1.
  • [35] Z. Liu, C. Zhao, I. Fedorov, B. Soran, D. Choudhary, R. Krishnamoorthi, V. Chandra, Y. Tian, and T. Blankevoort (2024) Spinquant: llm quantization with learned rotations. arXiv preprint arXiv:2405.16406. Cited by: TABLE X, TABLE IX, §5.
  • [36] Z. Liu, F. Li, G. Li, and J. Cheng (2021) EBERT: efficient bert inference with dynamic structured pruning. In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021, pp. 4814–4823. Cited by: §2.1.
  • [37] J. J. Park, P. Florence, J. Straub, R. Newcombe, and S. Lovegrove (2019) Deepsdf: learning continuous signed distance functions for shape representation. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pp. 165–174. Cited by: §2.2.
  • [38] S. Park, H. Choi, and U. Kang (2023) Accurate retraining-free pruning for pretrained encoder-based language models. In The Twelfth International Conference on Learning Representations, Cited by: §2.1.
  • [39] M. T. Pilehvar and J. Camacho-Collados (2019) WiC: the word-in-context dataset for evaluating context-sensitive meaning representations. In Proceedings of the 2019 conference of the North American chapter of the association for computational linguistics: Human language technologies, volume 1 (Long and short papers), pp. 1267–1273. Cited by: §5.
  • [40] P. Ram and K. Sinha (2019) Revisiting kd-tree for nearest neighbor search. In Proceedings of the 25th acm sigkdd international conference on knowledge discovery & data mining, pp. 1378–1388. Cited by: 1st item.
  • [41] M. Roemmele, C. A. Bejan, and A. S. Gordon (2011) Choice of plausible alternatives: an evaluation of commonsense causal reasoning.. In AAAI spring symposium: logical formalizations of commonsense reasoning, pp. 90–95. Cited by: §5.
  • [42] O. Ronneberger, P. Fischer, and T. Brox (2015) U-net: convolutional networks for biomedical image segmentation. External Links: 1505.04597 Cited by: §1, TABLE III, TABLE VIII, §5, TABLE XII.
  • [43] A. A. Rusu, D. Rao, J. Sygnowski, O. Vinyals, R. Pascanu, S. Osindero, and R. Hadsell (2018) Meta-learning with latent embedding optimization. In International Conference on Learning Representations, Cited by: §2.3.
  • [44] P. Sarlin, D. DeTone, T. Malisiewicz, and A. Rabinovich (2020) Superglue: learning feature matching with graph neural networks. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pp. 4938–4947. Cited by: §5.
  • [45] S. Shuvaev, D. Lachi, A. Koulakov, and A. Zador (2024) Encoding innate ability through a genomic bottleneck. Proceedings of the National Academy of Sciences 121 (38), pp. e2409160121. Cited by: §1, §3.1.
  • [46] K. O. Stanley, D. B. D’Ambrosio, and J. Gauci (2009) A hypercube-based encoding for evolving large-scale neural networks. Artificial life 15 (2), pp. 185–212. Cited by: §1, §3.1.
  • [47] M. Sun, Z. Liu, A. Bair, and J. Z. Kolter (2023) A simple and effective pruning approach for large language models. In The Twelfth International Conference on Learning Representations, Cited by: §2.1.
  • [48] Q. Team et al. (2024) Qwen2 technical report. arXiv preprint arXiv:2407.10671 2 (3). Cited by: TABLE III, TABLE V, TABLE V.
  • [49] T. Volk, E. Ben-David, O. Amosy, G. Chechik, and R. Reichart (2022) Example-based hypernetworks for out-of-distribution generalization. arXiv preprint arXiv:2203.14276. Cited by: §2.3.
  • [50] M. Xia, T. Gao, Z. Zeng, and D. Chen (2023) Sheared llama: accelerating language model pre-training via structured pruning. In The Twelfth International Conference on Learning Representations, Cited by: §2.1.
  • [51] M. Xia, T. Gao, Z. Zeng, and D. Chen (2024) Sheared llama: accelerating language model pre-training via structured pruning. External Links: 2310.06694 Cited by: TABLE III, TABLE V, TABLE V, §5, TABLE XIV.
  • [52] G. Xiao, J. Lin, M. Seznec, H. Wu, J. Demouth, and S. Han (2023) Smoothquant: accurate and efficient post-training quantization for large language models. In International Conference on Machine Learning, pp. 38087–38099. Cited by: §2.1, TABLE X, TABLE X, TABLE VI, TABLE IX, TABLE IX, §5.
  • [53] M. Xu, Y. L. Xu, and D. P. Mandic (2023) Tensorgpt: efficient compression of the embedding layer in llms based on the tensor-train decomposition. arXiv preprint arXiv:2307.00526. Cited by: §2.1.
  • [54] S. Xu, H. Li, B. Zhuang, J. Liu, J. Cao, C. Liang, and M. Tan (2020) Generative low-bitwidth data free quantization. In European conference on computer vision, pp. 1–17. Cited by: §5.
  • [55] J. Zhang, Y. Zhou, and R. Saab (2023) Post-training quantization for neural networks with provable guarantees. SIAM Journal on Mathematics of Data Science 5 (2), pp. 373–399. Cited by: §4.
  • [56] P. Zhang, G. Zeng, T. Wang, and W. Lu (2024) TinyLlama: an open-source small language model. External Links: 2401.02385 Cited by: TABLE III, TABLE V, TABLE V, TABLE XIV.
  • [57] S. Zhang, X. Liu, J. Liu, J. Gao, K. Duh, and B. Van Durme (2018) Record: bridging the gap between human and machine commonsense reading comprehension. arXiv preprint arXiv:1810.12885. Cited by: §5.
  • [58] S. Zhang, Z. Shen, and H. Yang (2022) Deep network approximation: achieving arbitrary accuracy with fixed number of neurons. Journal of Machine Learning Research 23 (276), pp. 1–60. Cited by: §3.3.
  • [59] S. Zhang, S. Roller, N. Goyal, M. Artetxe, M. Chen, S. Chen, C. Dewan, M. Diab, X. Li, X. V. Lin, et al. (2022) Opt: open pre-trained transformer language models. arXiv preprint arXiv:2205.01068. Cited by: TABLE XIV.
  • [60] Y. Zhao, C. Lin, K. Zhu, Z. Ye, L. Chen, S. Zheng, L. Ceze, A. Krishnamurthy, T. Chen, and B. Kasikci (2024) Atom: low-bit quantization for efficient and accurate llm serving. Proceedings of Machine Learning and Systems 6, pp. 196–209. Cited by: §2.1.
  • [61] K. Zhou, Q. Hou, R. Wang, and B. Guo (2008) Real-time kd-tree construction on graphics hardware. ACM Transactions on Graphics (TOG) 27 (5), pp. 1–11. Cited by: §1.
  • [62] X. Zhu, J. Li, Y. Liu, C. Ma, and W. Wang (2023) A survey on model compression for large language models. arXiv preprint arXiv:2308.07633. Cited by: §1, §2.1.
[Uncaptioned image] Fenglei Fan is currently an Assistant Professor with Department of Data Science, City University of Hong Kong. His primary research interests lie in NeuroAI and its applications in model compression and medical imaging. He was the recipient of the IBM AI Horizon Scholarship. His PhD dissertation was also selected as the recipient of the 2021 International Neural Network Society Doctoral Dissertation Award. His paper was selected as one of few 2024 CVPR Best Paper Award Candidates. He also won the IEEE Nuclear and Plasma Society IEEE TRPMS Best Paper Award.
[Uncaptioned image] Juntong Fan is a Ph.D. student in the Department of Data Science at City University of Hong Kong. He holds dual Master’s degrees: one in Mathematics from The Chinese University of Hong Kong and another in Social Computing from Xi’an Jiaotong-Liverpool University. His research primarily focuses on model compression theory and methodology. Prior to his doctoral studies, he served as a Research Assistant at CUHK from 2023 to 2024.
[Uncaptioned image] Dayang Wang is a visiting scholar at Department of Data Science, City University of Hong Kong. He earned his Ph.D. in Computer Engineering from the University of Massachusetts Lowell, where he was supervised by Prof. Hengyong Yu (IEEE Fellow). Prior to that, he received his bachelor’s degree in Computer Science from Beijing Normal University. His current research primarily focuses on the development of machine learning models and their applications in medical imaging.
[Uncaptioned image] Jingbo Zhang is a Master’s student at Northeastern University, China. Prior to this, he also finished his undergraduate study in Northeastern University in 2021. In 2024, he was invited to work as a Research Assistant at the Department of Mathematics, The Chinese University of Hong Kong. Now, he is a remote student intern at Department of Data Science, City University of Hong Kong. His research interest is model compression.
[Uncaptioned image] Zelin Dong is a Ph.D. student at the Department of Mathematics at the Chinese University of Hong Kong, having completed his undergraduate degree there in 2025. As a mentee in Prof. Fenglei Fan’s team (2022-2024), he conducted research in discrete computational geometry and received the UROP Gold Award in 2023. His research interests include partial differential equations and data science.
[Uncaptioned image] Shijun Zhang is an Assistant Professor at The Hong Kong Polytechnic University (PolyU). He graduated with a bachelor’s degree in mathematics from Wuhan University in 2016. Subsequently, he pursued his doctoral degree at National University of Singapore under the supervision of Professor Zuowei Shen and Professor Haizhao Yang, and successfully obtained his PhD in 2021. He used to work as a postdoctoral researcher at Duke University under the guidance of Professor Jianfeng Lu and Professor Hongkai Zhao. His primary research interest is in the (approximation) error analysis of neural networks.
[Uncaptioned image] Ge Wang (Fellow of IEEE, SPIE, AAPM, OSA, AIMBE, AAAS, and NAI) is the Clark & Crossan Chair Professor and Director of Biomedical Imaging Center at Rensselaer Polytechnic Institute (Troy, New York, USA). He pioneered the spiral cone-beam CT method in 1991 and published the first perspective on deep learning-based tomographic imaging in 2016. His honors include the IEEE EMBS Career Achievement Award, SPIE Meinel Technology Award, Sigma Xi Chubb Award for Innovation, RPI Wiley Distinguished Faculty Award, IEEE R1 Outstanding Teaching Award, and IEEE NPSS/NMISC Hoffman Medical Imaging Scientist Award.
[Uncaptioned image] Tieyong Zeng received the B.S. degree from Peking University, Beijing, China, in 2000, the M.S degree from Ecole Polytechnique, Palaiseau, France, in 2004, and the Ph.D. degree from the University of Paris XIII, Paris, France, in 2007. He is currently a Professor in the Department of Mathematics, The Chinese University of Hong Kong, Hong Kong. His research interests are image processing, machine learning, and scientific computing.